# Spectrality of Prime Size Tiles

### v3

@@ 1 Introduction
Here $p$ points being in general linear positions means they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$. It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [17, Corollary 2.4].
The proof of the first theorem applies several techniques: First we reduce the problem to $\mathbb{Z}_{n}^{d}$ by using periodicity, then we produce a size $p$ spectrum from $p$-subgroups of $\mathbb{Z}_{n}^{d}$. The construction involves dissecting the group into equivalence classes formed by generators of cyclic subgroups, while the existence of such a set being included in the zero set of $A$ is shown by contradiction. Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$, thus leading to a contradiction if $A$ does not admit such a spectrum.
@@ 2 Preliminaries
The way that Lemma 2 is stated here is slightly different from its original version in [20, Theorem 17], but it is clear that if $\{(t_{1},0,\ldots,0),\ldots,(0,\ldots,0,t_{d})\}$ is a set of periods of $B$, then the least common multiple of $t_{1},\ldots,t_{d}$ can be taken as $n$, and $B^{\prime}$ can be obtained by removing all duplications from $\pi_{n}(B)$.

### v4

@@ 1 Introduction
Here $p$ points being in general linear positions means they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$. It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [18, Corollary 2.4].
The proof of the first theorem applies several techniques: First we reduce the problem to $\mathbb{Z}_{n}^{d}$ by using periodicity, then we produce a size $p$ spectrum from $p$-subgroups of $\mathbb{Z}_{n}^{d}$. The construction involves dissecting the group into equivalence classes formed by generators of cyclic subgroups, while the existence of such a set being included in the zero set of $A$ is shown by contradiction. Projecting onto $\mathbb{Z}_{p^{k}}^{d}$ (where $p^{k}$ is the largest power of $p$ that divides $n$) and invoking the uncertainty principle, we are able to show that if the tiling complement annihilates all $p$-subgroups (which will be the case if $A$ annihilates none), then its size must be divisible by $p^{kd}$ (a statement that is similar to the Coven-Meyerowitz T1 property in [2]), thus leading to a contradiction if $A$ does not admit such a spectrum.
@@ 2 Preliminaries
The way that Lemma 2 is stated here is slightly different from its original version in [21, Theorem 17], but it is clear that if $\{(t_{1},0,\ldots,0),\ldots,(0,\ldots,0,t_{d})\}$ is a set of periods of $B$, then the least common multiple of $t_{1},\ldots,t_{d}$ can be taken as $n$, and $B^{\prime}$ can be obtained by removing all duplications from $\pi_{n}(B)$.